OpenStudy (macabbey101):

Solve and express the solution set in simplest form. 4x-1/3 = 7/1 A. {4/3} B. {7/3} C. {3/2} D. {11/2}

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 4x-\frac{1}{3}=7 }\) \(\LARGE\color{blue}{ \bf 4x=7+ \frac{1}{3}}\) \(\LARGE\color{blue}{ \bf 4x=\frac{22}{3}}\) divide both sides by 4 ... \(\LARGE\color{blue}{ \bf 4x=\frac{22}{12} = \frac{11}{6} = 1\frac{5}{6}}\)

OpenStudy (macabbey101):

divide what by 4? @SolomonZelman

OpenStudy (solomonzelman):

the last step... \(\Huge\color{blue}{ \sf 4x= \frac{22 }{3} }\) \(\Huge\color{blue}{ \sf 4x÷4= \frac{22 }{12} ÷4}\) \(\Huge\color{blue}{ \sf x= \frac{22 }{12} }\)

OpenStudy (macabbey101):

im so confused. i got 11/24 for dividing 22/12 by 4. :(

OpenStudy (macabbey101):

@SolomonZelman

OpenStudy (macabbey101):

if you tell me the answer i will give you a medal!! please im in a hurry :( @SolomonZelman

OpenStudy (solomonzelman):

it is not any of the options you posted. it is 1 5/6

OpenStudy (anonymous):

Macabbey, if you're still confused, I'd be happy to help if Solomon allows it.